# Time Travel in LLMs: Tracing Data Contamination in Large Language Models

### v1

@@ Abstract
Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a potential major issue in understanding LLMs’ effectiveness on other tasks. We propose a straightforward yet effective method for identifying data contamination within LLMs. At its core, our approach starts by identifying potential contamination in individual instances that are drawn from a small random sample; using this information, our approach then assesses if an entire dataset partition is contaminated. To estimate contamination of individual instances, we employ “guided instruction:” a prompt consisting of the dataset name, partition type, and the initial segment of a reference instance, asking the LLM to complete it. An instance is flagged as contaminated if the LLM’s output either exactly or closely matches the latter segment of the reference. To understand if an entire partition is contaminated, we propose two ideas. The first idea marks a dataset partition as contaminated if the average overlap score with the reference instances (as measured by ROUGE or BLEURT) is statistically significantly better with the guided instruction vs. a general instruction that does not include the dataset and partition name. The second idea marks a dataset as contaminated if a classifier based on GPT-4 with in-context learning prompting marks multiple instances as contaminated. Our best method achieves an accuracy between 92% and 100% in detecting if an LLM is contaminated with seven datasets, containing train and test/validation partitions, when contrasted with manual evaluation by human expert. Further, our findings indicate that GPT-4 is contaminated with AG News, WNLI, and XSum datasets.
@@ 1 Introduction
The rise of Transformer networks (Vaswani et al. 2017) has spurred the development of large language models (LLMs), marking a new epoch in Natural Language Processing (NLP). This shift has led to an extensive range of LLMs (Touvron et al. 2023a; b; Biderman et al. 2023; Köpf et al. 2023; Chung et al. 2022; Penedo et al. 2023, inter-alia) which excel in various professional and academic benchmarks (Bang et al. 2023; Bubeck et al. 2023). Their superior performance is primarily attributed to the massive web data consumed by these billion/trillion-parameter LLMs during training. However, the impressive LLM performance observed on many downstream tasks (e.g., summarization, natural language inference, text classification) may be inflated due to data contamination, i.e., the presence of test data from these downstream tasks in the training data of LLMs. Guaranteeing lack of contamination is not trivial due to two potential sources of contamination: directly from ingesting the official version of a dataset (easier to control), and indirectly through duplicated data found somewhere on the web (nearly impossible to control). [^1] While dataset licensing reduces indirect contamination to a certain extent, it does not eliminate it. For example, websites such as the Hugging Face page for datasets (Wolf et al. 2020) currently host copies of the OntoNotes (Weischedel et al. 2013) and CoNLL-2003 (Tjong Kim Sang and De Meulder 2003) datasets, despite the fact that their respective licenses prohibit it. The potential of data contamination is especially relevant for closed models such as the GPT-3/3.5 family (Brown et al. 2020) and GPT-4 (OpenAI 2023; Bubeck et al. 2023), and, needless to say, raises questions on the validity of evaluations conducted so far (Chang et al. 2023; Zhu et al. 2023; Bordt and von Luxburg 2023; Ray 2023; Penedo et al. 2023).
@@
To address this issue, we propose an inexpensive and robust approach to automatically detect data contamination for a given dataset partition. Intuitively, our method starts by identifying potential contamination in individual instances that are drawn from a small random sample of the corresponding dataset partition (we used samples of 10 instances in this work). Using this information, our approach then assesses if an entire dataset partition is contaminated.
@@
More formally, to identify contamination of individual instances we employ a “guided instruction:” an LLM prompt that integrates distinct identifiers from the source dataset from which the reference instance originates. Such information includes the dataset name, its partition (e.g., train, test, or validation), and a randomly selected portion of the reference instance, complemented by its label when relevant. With these signals in the prompt, we instruct the LLM to finish the given partial instance. To understand if an entire dataset partition is contaminated, we propose two heuristics. The first heuristic states that a partition is likely to be contaminated if the average overlap score between generated and reference instances (as measured by ROUGE-L (Lin 2004) or BLEURT (Sellam, Das, and Parikh 2020)) observed with the guided instruction, is statistically significantly larger than the one measured with a general instruction, which does not include the dataset and partition name. The second heuristic labels a partition as contaminated if a classifier based on GPT-4 with in-context learning (ICL; Brown et al. (2020)) marks at least one instance as an exact match with the reference or at least two instances as near-exact matches, where near-exact matches indicate instances that exhibit considerable semantic and lexical alignment with the reference instance.
@@
The primary contributions of this paper are as follows: (1) We propose a novel data contamination detection method for LLMs that is inexpensive and robust. As indicated above, our method combines a “guided instruction” to complete instances randomly drawn from the investigated dataset and several heuristics to generalize from instance- to partition-level contamination decisions.
@@
(2) We evaluate our proposed method in 28 distinct scenarios, using two state-of-the-art LLMs—GPT-4 and GPT-3.5—seven different datasets employed in classification, summarization, and natural language inference (NLI) tasks. For each dataset, we investigate potential data contamination in both the train and test splits (or validation, if the test partition is not publicly available with its instance labels). Our evaluation indicates that our best method is the one that combines the guided instruction to complete individual instances with the GPT-4 ICL classifier. This method obtains an accuracy that ranges between 92% and 100% for the two language models when compared with dataset contamination labels produced by a human domain expert.
@@
(3) Our analysis indicates that GPT-4 showed evidence of contamination within the test segments for AG News (Zhang, Zhao, and LeCun 2015), WNLI (Wang et al. 2018), and XSum (Narayan, Cohen, and Lapata 2018) datasets. This finding supports the observation that data contamination is a serious issue that must be considered in downstream evaluations using LLMs.
@@ 2 Related Work
Despite its importance, the topic of data contamination is not as thoroughly examined as its closely related field, data memorization (Carlini et al. 2023; Kandpal, Wallace, and Raffel 2022; Carlini et al. 2021; Razeghi et al. 2022). Among the limited investigations focusing specifically on data contamination in LLMs, we find notable examples in Radford et al. (2019) and Brown et al. (2020) on GPT-2 and GPT-3, respectively. They used high-order $n$-grams (e.g., 13-gram) to detect overlapping content between the training data and the evaluated dataset. Most research subsequent to Brown et al. (2020) adopted similar methods for detecting data contamination (Touvron et al. 2023b; Du et al. 2022; Chowdhery et al. 2022; Wei et al. 2021). However, the scope of existing research has been predominantly confined to model providers, and it encounters specific limitations, particularly when applied to closed-source LLMs. These limitations primarily involve the need for access to pretraining data (Brown et al. 2020; Du et al. 2022; Wei et al. 2021), the requirement for substantial computational resources (Touvron et al. 2023b), or the need for extensive manual labor (Chowdhery et al. 2022). Our approach aims to overcome these hurdles, enabling the assessment of data contamination in scenarios where the training data is either not openly accessible or when significant computational hardware is not available despite having access to the training data.
The work of Sainz et al. (2023) represents the most pertinent effort to tackle data contamination in LLMs. This effort prompted ChatGPT (also known as GPT-3.5-turbo) to generate initial instances from the different splits of a dataset. The underlying assumption here is that if an LLM can reproduce these instances, it must have been trained using that particular split. However, while this approach is prompt-sensitive, our research shows that this method can often prove unreliable and subject to failure due to the sparsity introduced by the request to reproduce the initial instances of the dataset split. Throughout this paper, we refer to this approach as “Did-ChatGPT-Cheat?,” taking inspiration from the title of the referenced blog post.
In this paper, we propose a novel technique for detecting data contamination in LLMs. In a nutshell, our approach instructs the LLM to produce the second segment of a reference instance when its initial part is presented in the input prompt. We subsequently assess the output for exact or nearly exact matches. To extrapolate contamination across the entire dataset partition, we use evaluation measures such as ROUGE, BLEURT, GPT-4 evaluation, and human judgment. When tested on seven datasets using GPT-4 and GPT-3.5, our method provides the trace of leaked dataset splits.
@@ 3 Approach
In our approach, we operate under a core assumption that we lack direct access to the training data of the LLMs. Given this premise, our detection strategy for data contamination is anchored by two pivotal insights. First, we examine individual instances within a dataset partition to spot contamination at the instance level. Second, given that LLMs are trained on grand-scale data, detecting contaminated instances can act as a signal of broader contamination. As a result, the associated partition can be labeled as being leaked to the LLM’s training data.
To discern contamination at the instance level, we focus on replicating instances by the LLM. In this context, exact replicas of instances serve as red flags for contamination in the corresponding partition, confirming these instances were part of the LLM’s training data. However, due to the inherent probabilistic behavior of LLMs, achieving perfect replicas is not always possible. Nevertheless, instances that are closely replicated have a twofold function: while they can offer insightful indications of potential contamination, the fact that many datasets draw from web-based sources implies that partial replicas can also arise by happenstance. This overlap introduces uncertainty in drawing definitive conclusions about the underlying partition. Thus, it is essential to check for consistent and significant signs of contamination within the partition.
In the following sections, we first elaborate on our method and the necessary components to compel LLM into reproducing dataset instances. We then delve into the procedure for evaluating the contamination status of an entire partition based on these instances.
@@ 3.1 Detecting Instance-level Contamination
Our procedure identifies contamination at the instance level in four steps. First, we steer the LLM towards the (potentially-contaminated) dataset using a guided instruction that integrates the dataset name and the partition of interest. Next, to detect potential data contamination, we provide an initial random segment of an instance and its label if it is available. The LLM is then instructed to complete it (see below for the instructions used). Lastly, we use a series of methods to assess how closely the LLM’s output, under guidance, matches the subsequent portion of the reference instance. An example of a guided versus a general instruction is depicted in Figure 1.
(1) Guided Instruction—A Means to Navigate the LLM’s Domain. By employing instruction-tuning on top of causal language modeling (CLM; Vaswani et al. (2017); Radford et al. (2018)), LLMs can be guided by human directives (Wei et al. 2022; Sanh et al. 2022; Chung et al. 2022). This serves as a tool for controlling the LLM’s domain using natural language. Thus, we form a guided instruction that incorporates the dataset and split names in the prompt, thereby directing the LLM towards the underlying dataset. A comprehensive list of all the instructions used in this study for different tasks/datasets can be found in Table 3 in Appendix A.
(2) Unraveling Data History with Causal Language Modeling. Primarily, data contamination occurs during the CLM phase since it constitutes the largest part of training in LLMs and utilizes web data. Without instruction tuning, an LLM only attempts to complete an input prompt based on data seen during the CLM training phase (Ouyang et al. 2022). Notable models that exhibit this behavior include GPT-2, GPT-3, and the LLaMA models (Touvron et al. 2023a; b). We, therefore, employ the next token prediction mechanism to trace data history. In particular, we feed the model the initial segment of a random dataset instance from a particular split, prompting it to finish the instance. For labeled instances, we integrate the corresponding labels in the input prompt. This reflects that if an instance was ingested during the LLM’s training phase, its label was ingested too.
Therefore, once the LLM is prompted with guided instruction, its output mirrors the subsequent segment of the reference instance, resulting in either an exact, near-exact, or inexact match. The criteria we use to categorize these matches are detailed in step 4.
(3) General Instruction—An Alternate Facet of Causal Language Modeling. We also formulate a general instruction to measure the impact of the guidance given in the guided instruction. This general instruction only requests the completion of the partial input instance without specifying the dataset or its partition. As a result, when using this general instruction, the generated sequence solely relies on the CLM training phase, akin to autoregressive models without instruction tuning. This enables us to establish a baseline for generated random replicas and assess how much the guided instruction influences the LLM-generated part of the input instance. We assess this influence in terms of overlap, semantics, and structural similarity with the reference instance. This analysis is crucial as even when the output of LLM does not perfectly match the reference instance, it still enables us to detect potential signs of contamination.
(4) Measuring Instance-level Contamination.
BLEURT & ROUGE: To quantify the overlap between a generated and a reference instance we employ ROUGE (Lin 2004) and BLEURT (Sellam, Das, and Parikh 2020) scores. While ROUGE assesses lexical similarity, BLEURT gauges the semantic relevance and fluency of the resulting sequence with respect to the reference instance.
GPT-4 Evaluation: While both BLEURT and ROUGE quantify the overlap between the generated and reference instances, they fall short of pinpointing near-exact matches. To bridge this gap, we adopt ICL prompting (Brown et al. 2020)) to dictate the detection of exact/near-exact matches based on human judgment. Specifically, this method entails presenting a few representative examples of exact and near-exact matches—sourced from human evaluation—within the input prompt and then proceeding to assess all other instances. We choose GPT-4 for this task as it requires no specialized prompting technique (Bubeck et al. 2023), enhancing the reliability of its results. A visual representation of the ICL prompt used in our study can be seen in Figure 2 (Appendix C). Also, detailed examples, including their ROUGE and BLEURT scores accompanied by both human and GPT-4 evaluations, are provided in Table 4 (Appendix D).

### v2

@@ Abstract
Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a potential major issue in measuring LLMs’ real effectiveness on other tasks. We propose a straightforward yet effective method for identifying data contamination within LLMs. At its core, our approach starts by identifying potential contamination at the instance level; using this information, our approach then assesses wider contamination at the partition level. To estimate contamination of individual instances, we employ “guided instruction:” a prompt consisting of the dataset name, partition type, and the random-length initial segment of a reference instance, asking the LLM to complete it. An instance is flagged as contaminated if the LLM’s output either exactly or nearly matches the latter segment of the reference. To understand if an entire partition is contaminated, we propose two ideas. The first idea marks a dataset partition as contaminated if the average overlap score with the reference instances (as measured by ROUGE-L or BLEURT) is statistically significantly better with the completions from guided instruction compared to a “general instruction” that does not include the dataset and partition name. The second idea marks a dataset partition as contaminated if a classifier based on GPT-4 with few-shot in-context learning prompt marks multiple generated completions as exact/near-exact matches of the corresponding reference instances. Our best method achieves an accuracy between 92% and 100% in detecting if an LLM is contaminated with seven datasets, containing train and test/validation partitions, when contrasted with manual evaluation by human experts. Further, our findings indicate that GPT-4 is contaminated with AG News, WNLI, and XSum datasets.
@@ 1 Introduction
The rise of Transformer networks (Vaswani et al. 2017) has spurred the development of large language models (LLMs), marking a new epoch in Natural Language Processing (NLP). This shift has led to an extensive range of LLMs (Touvron et al. 2023a; b; Biderman et al. 2023; Köpf et al. 2023; Chung et al. 2022; Penedo et al. 2023, inter-alia) which excel in various professional and academic benchmarks (Bang et al. 2023; Bubeck et al. 2023). Their superior performance is primarily attributed to the massive web data consumed by these billion/trillion-parameter LLMs during training. However, the impressive LLM performance observed on many downstream tasks (e.g., summarization, natural language inference, text classification) may be inflated due to data contamination, i.e., the presence of test data from these downstream tasks in the pre-training data of LLMs. Guaranteeing lack of contamination is not trivial due to two potential sources of contamination: directly from ingesting the official version of a dataset (easier to control), and indirectly through duplicated data found somewhere on the web (nearly impossible to control). [^1] While dataset licensing reduces indirect contamination to a certain extent, it does not eliminate it. For example, websites such as the Hugging Face page for datasets (Wolf et al. 2020) currently host copies of the OntoNotes (Weischedel et al. 2013) and CoNLL-2003 (Tjong Kim Sang & De Meulder 2003) datasets, despite the fact that their respective licenses prohibit it. The potential of data contamination is especially relevant for closed models such as the GPT-3/3.5 family (Brown et al. 2020) and GPT-4 (OpenAI 2023; Bubeck et al. 2023), and, needless to say, raises questions on the validity of evaluations and benchmarks conducted so far (Chang et al. 2023; Zhu et al. 2023; Bordt & von Luxburg 2023; Ray 2023; Penedo et al. 2023).
@@
To address this issue, we propose an inexpensive and robust approach to detect data contamination for a given dataset partition automatically. Importantly, our approach functions under two realistic assumptions: (a) we lack direct access to the pre-training data of the LLMs, and (b) we have limited computational resources. Intuitively, our method starts by identifying potential contamination in individual instances that are drawn from a small random sample of the corresponding dataset partition (we use samples of 10 instances in this work). Using the information obtains from individual instances, our approach then assesses if an entire dataset partition is contaminated.
@@
More formally, to identify contamination of individual instances, we employ a “guided instruction:” a prompt that integrates distinct identifiers from the source dataset from which the reference instance originates. Such information includes the dataset name, its partition (e.g., train, test, or validation), and a randomly selected initial portion of the reference instance, complemented by its label when relevant. With these signals in the prompt, we instruct the LLM to finish the given partial instance. Using these generated individual completions, we propose two heuristics to estimate if an entire dataset partition is contaminated. The first heuristic states that a partition is likely to be contaminated if the average overlap score between generated completions and reference instances (as measured by ROUGE-L (Lin 2004) and BLEURT (Sellam et al. 2020)) observed with the guided instruction is statistically significantly larger than the one measured with a “general instruction,” which does not include the dataset and partition name. The second heuristic labels a partition as contaminated if a classifier based on GPT-4 with few-shot in-context learning (ICL; Brown et al. (2020)) marks at least one generated completion as an exact match with the reference instance or at least two generated completions as near-exact matches, where near-exact match indicates a completion that exhibits considerable semantic and lexical alignment with the reference instance.
The primary contributions of this paper are as follows:
@@
(1) We propose a novel data contamination detection method for LLMs that is inexpensive and robust. As indicated above, our method combines a “guided instruction” to complete partial instances randomly drawn from the investigated dataset partition and several heuristics to generalize from instance- to partition-level contamination decisions.
@@
(2) We evaluate our proposed methods in 28 distinct scenarios. These scenarios are created by two state-of-the-art LLMs: GPT-3.5 and GPT-4, and span seven datasets for classification, summarization, and natural language inference (NLI) tasks. The rationale behind the 28 scenarios is that for each dataset, we separately explore potential data contamination in the train and test splits (or the validation set, in cases where the labeled test set is not publicly available). Our evaluation indicates that our best method is the one that uses guided instruction to complete partial instances, and the one that evaluates these completions by the GPT-4 few-shot ICL classifier, achieving 92%–100% accuracy compared to contamination labels assigned by human experts for dataset partitions.
@@
(3) Our analysis indicates that GPT-4 showed evidence of contamination within the test partitions of AG News (Zhang et al. 2015), WNLI (Wang et al. 2018), and XSum (Narayan et al. 2018) datasets. These findings support the observation that data contamination is a serious issue that must be considered in downstream evaluations when using LLMs.
@@ 2 Related Work
Despite its importance, the topic of data contamination is not as thoroughly examined as its closely related field, data memorization (Carlini et al. 2023; Kandpal et al. 2022; Carlini et al. 2021; Razeghi et al. 2022). Among the limited investigations focusing specifically on data contamination in LLMs, we find notable examples in Radford et al. (2019) and Brown et al. (2020) on GPT-2 and GPT-3, respectively. They used high-order $n$-grams (e.g., 13-gram) to detect overlapping content between the pre-training data and the evaluation dataset. Most research subsequent to Brown et al. (2020) adopted similar methods for detecting data contamination (Touvron et al. 2023b; Du et al. 2022; Chowdhery et al. 2022; Wei et al. 2021), and most recently, substring matching for GPT-4 (OpenAI 2023). However, the scope of existing research has been predominantly confined to model providers, and it encounters specific limitations, particularly when applied to closed-source LLMs. These limitations primarily involve the need for access to pre-training data (Brown et al. 2020; Du et al. 2022; Wei et al. 2021), the requirement for substantial computational resources (Touvron et al. 2023b), or the need for extensive manual labor (Chowdhery et al. 2022). Our approach aims to overcome these hurdles, enabling the assessment of data contamination in scenarios where the pre-training data is either not openly accessible or when significant computational hardware is not available despite having access to the pre-training data.
Our paper is closest in spirit to the work of Sainz et al. (2023), who also detected contamination when access to the pre-training data is not available. This effort prompted ChatGPT, particularly when GPT-3.5 is its base model, to generate the first instances from different dataset partitions. The underlying assumption here is that if an LLM can reproduce dataset instances, it must have been trained using that particular split. However, our research shows that this method can be unreliable and subject to failure. Such failures can result either from the sparsity introduced by the request to reproduce the first instances of a dataset split or from the inability to bypass the safety filters set by the model provider when the model is asked to generate copyrighted content like dataset instances. Throughout this paper, we refer to this approach as “ChatGPT-Cheat?,” taking inspiration from the title of the referenced blog post.
@@ 3 Approach
In our approach, we operate under two core assumptions: (1) lacking direct access to the pre-training data of the LLMs, and (2) having limited computational resources. Given these premises, our detection strategy for data contamination is anchored by two pivotal insights. First, we examine individual instances within a dataset partition to spot contamination at the instance level. Second, given that LLMs are pre-trained on large-scale data, detecting contaminated instances can act as a signal of broader contamination. As a result, the associated partition can be labeled as being leaked to the LLM’s pre-training data.
To discern contamination at the instance level, we focus on replicating instances by the LLM. In this context, exact replicas of instances serve as red flags for contamination in the corresponding partition. Note that, due to the inherent probabilistic behavior of LLMs, achieving perfect replicas is not always possible even when contamination is certain. Nevertheless, instances that are closely replicated have a twofold function: while they can offer insightful indications of potential contamination, the fact that many datasets draw from web-based sources implies that partial replicas can also arise by happenstance. This overlap introduces uncertainty in drawing a definitive conclusion about the underlying partition. Thus, it is essential to check for consistent and significant signs of contamination within the partition.
In the following sections, we first elaborate on our method and the necessary components to compel LLM into reproducing dataset instances. We then delve into the procedure for evaluating the contamination status of existing LLMs for an entire partition based on these instances. Furthermore, leveraging the fine-tuning option offered by OpenAI for the GPT-3.5 base model, we undertake a study in which we intentionally contaminate the GPT-3.5 base model with partitions that our method detected as uncontaminated. Subsequently, we subject the contaminated GPT-3.5 to our technique, further showcasing our method’s effectiveness in pinpointing data contamination within LLMs.
@@ 3.1 Detecting Instance-level Contamination
3.1.1 Components to Measure Instance-level Contamination
To gauge instance-level contamination, we utilize two distinct methods: the first leverages BLEURT and ROUGE-L scores, while the second draws on few-shot ICL prompting with GPT-4. While each of these methods requires specific components to be employed, the first two components are shared between both. The third component, the general instruction, is exclusive to the first method. For both methods, we begin our process by steering the LLM towards the (potentially contaminated) dataset partition using guided instruction that integrates the dataset name and the partition of interest. Next, we provide the random-length initial segment of a randomly selected instance and its label if it is available. The LLM is then instructed to complete it. For the first method, we repeat this step using general instruction that omits the dataset and partition name. An example of a guided versus a general instruction is depicted in Figure 1. We detail all the required components below.
@@ 3.1.1 Components to Measure Instance-level Contamination
(1) Guided Instruction—A Means to Navigate the LLM’s Domain. By employing instruction-tuning on top of causal language modeling (CLM; Vaswani et al. (2017); Radford et al. (2018)), LLMs can be guided by human directives (Wei et al. 2022; Sanh et al. 2022; Chung et al. 2022). This serves as a tool for controlling the LLM’s domain using natural language. Thus, we form guided instruction such that it incorporates the dataset and split name in the input prompt, thereby directing the LLM towards the underlying dataset split. A comprehensive list of all the instructions used in this study for different tasks/datasets can be found in Table 5 in Appendix A.
(2) Unraveling Data History with Causal Language Modeling. Primarily, data contamination occurs during the CLM pre-training phase since it constitutes the largest part of training in LLMs and utilizes web data. Without instruction tuning, an LLM only attempts to complete an input prompt based on data seen during the CLM pre-training phase (Ouyang et al. 2022). Notable models that exhibit this behavior include GPT-2 and GPT-3. We, therefore, employ the next token prediction mechanism to trace data history. In particular, we feed the model the variable-length initial segment of a dataset instance, chosen randomly from a particular split, prompting it to finish the partial instance. For labeled instances, we integrate the corresponding labels in the input prompt. This reflects that if an instance was ingested during the LLM’s pre-training, its label was ingested too.
Therefore, once a contaminated LLM is prompted with guided instruction, its output should mirror the subsequent segment of the reference instance under the guidance of the dataset and split name.
(3) General Instruction—An Alternative Facet of Causal Language Modeling. We formulate the general instruction to measure the impact of the guidance given in the guided instruction. This general instruction only requests the completion of the partial instance without specifying the dataset or its partition. As a result, when using this instruction, the generated sequence solely relies on the CLM pre-training phase, akin to autoregressive models without instruction tuning. This enables us to establish a baseline for generated random replicas and assess how much the guided instruction influences the LLM-generated part of the input partial instance. We assess this influence in terms of overlap, semantics, and structural similarity with the reference instance. This analysis is crucial as even when the output of LLM does not perfectly match the reference instance, it still enables us to detect potential signs of contamination.
3.1.2 Measuring Instance-level Contamination
We introduce two methods for measuring contamination at the instance level:
BLEURT & ROUGE-L: To quantify the overlap between the completions—produced under both guided and general instructions—and reference instances, we employ two metrics: ROUGE-L (Lin 2004) and BLEURT (Sellam et al. 2020). While ROUGE-L assesses lexical similarity, BLEURT gauges the semantic relevance and fluency of the resulting sequence with respect to the reference instance. Instance-level contamination is detected if the average overlap scores from either metric, when applied to completions from the guided instruction, exceed those from the general instruction.
GPT-4 Evaluation: While both BLEURT and ROUGE-L quantify the overlap between the generated and reference instances, they fall short of pinpointing near-exact matches. To bridge this gap, we adopt few-shot ICL prompting (Brown et al. 2020) to dictate the detection of exact/near-exact matches based on human judgments (see Section  4: Human Evaluation for our definition of a near-exact match). Specifically, this method includes a few representative examples of exact and near-exact matches—sourced from human evaluations—in the prompt, which are used to assess all other generated completions. We chose GPT-4 for this task as it requires no specialized prompting technique (Bubeck et al. 2023), enhancing the reliability of its results. A visual representation of the few-shot ICL prompt used in our study can be seen in Figure 3 in Appendix B. Also, detailed examples, including their ROUGE-L and BLEURT scores, as well as both human and GPT-4 few-shot ICL evaluations, are listed in Table 6 in Appendix C.